OpenStudy (ervinloriaga):

A metal hollow bar whose cross section and dimension are shown below weighs 8x10^3 kg/m^3 and measure 2m in length ..determine the mass of the metal bar with the square hole section. Edge = 40cm -- square Edge = 60 cm -- hexagon

OpenStudy (mathmate):

I assume the edge measurements are exterior measurements. (it is normal to do exterior measurements. However, without the figure, we do not know the thickness (or any other undetermined information). So please post the figure or make a complete verbal description of information. By the way, the following might help you help yourself: (e=side length, L=longitudenal length) Following shapes are SOLID. shape cross section area volume Square e^2 Le^2 Regular Hexagon \(\Large \frac{3\sqrt{3}e}{2}\) \(\Large \frac{3\sqrt{3}eL}{2}\)

OpenStudy (sshayer):

after finding Volume ,multiply unit weight by Volume.